OpenStudy (anonymous):

Can anyone explain how to prove(sin x)(tan x cos x - cot x cos x) = 1 - 2 cos2x This proving trig functions has got me confused

Nnesha (nnesha):

write tan in terms of sin over cos \[\huge\rm tan \theta = \frac{ \sin }{ \cos }\] change that and remember cot is reciprocal of tan

Nnesha (nnesha):

\[\large\rm sinx (\color{reD}{tanx} cosx - \color{red}{cotx} cosx)\]\[\large\rm sinx(\color{Red}{\frac{ \sin x}{ cosx }} cosx - \cot x cosx)\] cot x = what ?

OpenStudy (anonymous):

Give me a sec

OpenStudy (pawanyadav):

Sinx(sinx/cosx×cos-cosx/sinx×cosx) =sin^2x-cos^2x. (Multiply sin inside the bracket) =-cos2x

OpenStudy (pawanyadav):

The answer is incorrect

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @lala102 Give me a sec \(\color{blue}{\text{End of Quote}}\) okay :-)

OpenStudy (anonymous):

Okay i am so confused gahhhh!!! I really hate trig. Do I distribute? I do not know what I am doing, but I'm trying. Math is just my enemy

Nnesha (nnesha):

where is your work ??

Nnesha (nnesha):

cot x = what ? if tan = sin/cos

OpenStudy (pawanyadav):

Want to solve it then love it play with it . It's not your enemy.

OpenStudy (anonymous):

Oh okay I got you now cotx is 1/tanx right

Nnesha (nnesha):

nope tan and cot are reciprocal of each other so if tan = sin /cos then cot = cos /sin

Nnesha (nnesha):

\[\large\rm sinx(\color{Red}{\frac{ \sin x}{ cosx }} cosx - \frac{cosx}{sin} cosx)\] now solve the parentheses first

OpenStudy (anonymous):

Um okay.... Man I must be looking at the wrong set of notes. I'm sorry

Nnesha (nnesha):

what do yo mean ??

Nnesha (nnesha):

both sides are equal you can verify

OpenStudy (anonymous):

I'm sorry I'm going slow its somethinwrong with my phone

Nnesha (nnesha):

alright let sin x = a and cos x = b \[\large\rm a(\color{Red}{\frac{ a}{ b }}( b)- \frac{b}{a} (b))\] simple algebra can you solve it now

OpenStudy (anonymous):

Yes thank you!

Nnesha (nnesha):

got it ?